# The Perception of Phase Intercept Distortion and its Application in Data Augmentation

### v1

@@ 1.1 Phase-intercept Distortion
$x(t)=\sin(\omega_{0}t+\phi)=\frac{1}{2i}(e^{i\omega_{0}t}e^{i\phi}-e^{-i\omega_{0}t}e^{-i\phi})$
Note that we must add the positive phase to the positive frequencies and the negative of that phase to the negative frequencies. This can be verified by using a single cosine tone and the effect of a shift in phase in its frequency domain.
$\hat{x}(\omega)=\frac{1}{2i}[\delta(\omega-\omega_{0})e^{i\phi}-\delta(\omega+\omega_{0})e^{-i\phi}]$
@@
A useful property of the Hilbert transform of a real-valued signal $x(t):\mathbb{R}\rightarrow\mathbb{\mathbb{R}}$, is that it is orthogonal to the original signal, as can be easily proved by showing that its inner product with $\mathcal{H}(x)(t)$ is equal to zero. This property makes it fundamental in applications such as analytic signal representation [14, 15], envelope detection [16], phase retrieval [17], single-sideband (SSB) modulation, and instantaneous frequency estimation [18].
@@
This can be shown by expanding 4 and rearranging the terms. In summary, phase-intercept distortion can be modeled by constructing the analytic representation of the signal using the Hilbert transform, rotating it, and extracting the real part of this complex signal.
@@ 1.2 Perception of Phase-intercept Distortion
The perception of phase-intercept distortion has not been explored as much as other common forms of phase distortion. One specific case that has been studied is polarity reversal—corresponding to a phase shift where $\theta=\pm 180^{\circ}$ in 7. Though most researchers (and audio engineers) assume that polarity reversal is inaudible, which is consistent with experimental evidence [19], there have also been some claims that it is audible [7, 8]. The Wood Effect [^1] Named after Charles L. Wood. [8] is one dramatic edge case where polarity reversal creates a minute difference in timbre for lower frequencies. This effect occurs when inverting a sine wave which has been clipped on only one half of each cycle; The resulting polarity-inverted signals sound slightly different, showing that our ears do treat the clipped portion differently when it is presented as a compression or as a rarefaction. No other cases of perceivable phase inversion have been reported.
@@ 2.1 Participants
Forty-six participants were recruited via different professional and university mailing lists.  [^2] This experiment was conducted with the approval of the Georgia Tech’s Institution Review Board (IRB) (ethics board). Participants were only required to be physically present in the United States, at least 18 years old, and have no history of hearing loss (screened through self-report). Only 25 of our 46 participants completed 100% of the survey and we discarded data from one participant who reported a hearing disability, leaving complete data from 24 participants.
We anticipated that individual differences between participants (age, gender, music background, etc.) might affect their ability to distinguish phase-intercept distortion. Thus, through a pre-questionnaire, we collected information about the participants’ age, gender, and experience with music, musical instruments, music production, mixing engineering, and high-fidelity listening habits. Twenty-one participants identified as male (mean age: 29.3 years) and the rest as female (mean age: 38.3 years); six of the participants were above the age of 35. Fourteen participants identified as musicians. In the final analysis, none of these factors seem to affect participants’ performance.
@@ 2.3 Audio Editing and Manipulation
We computed the Hilbert transform of each of the thirty stimuli to construct the analytic signals using 6. To finally apply phase-intercept distortion, the sampled $\theta$ values were then inputted into 7.
@@ 3 Results
For each trial, the participant’s response (A or B) can be coded correct or incorrect. If our hypothesis is true, participants’ true success rate should be approximately 50%. On average, our participants selected correctly in 49.44% of trials (Table 1). A one-sample t-test was performed to check if this accuracy was significantly different from the chance level of 50%, and the result was not significant. However, a non-significant test result might arise from a small sample size, especially if the true effect is small. We thus, focus on Bayesian statistical analysis, which is more appropriate for this analysis.
Table 1: Perceptual Test Results
@@
Mean 51.67% 44.58% 52.08% 49.44%
Median 50% 45% 50% 50%
Standard deviation 15.18% 19.36% 16.07% 8.59%
t-statistic 0.526 -1.342 0.622 -0.310
p-value 0.604 0.193 0.540 0.759
Samples 24
Figure 3: Distribution of response accuracy of 24 participants.
@@
Out of the 720 total questions answered by the participants, they successfully selected the undistorted stimuli 356 times and failed 364 times. Thus, the Bayesian posterior distribution for $q$, given the naive prior distribution we started with, is $q_{posterior}\sim\textnormal{Beta}(\alpha=1+356,\ \beta=1+364)$ (Fig. 4). As can be seen in Fig. 4, even if we begin with flat prior belief about $q$, the data from this experiment leads to a posterior belief about $q$ that is tightly centered around $q=0.5$, with 95% of this posterior distribution falling between 0.458 and 0.531. These results are consistent with the experimental hypothesis that phase-intercept distortion has no perceptible effects.
@@
Table 1 reports the results for each of the categories: music, speech and other. The median results are 50% for all categories except speech, which has a median of 45%. The mean scores are similar for music and other, while the mean score of speech is worse than 50%.
@@
To ensure that participants’ fatiguing over the course of the survey was not a confound, Fig. 5 plots the question-wise scores, with a fitted trend line. This least-squares regression line has a negligible slope of $-0.002$. This suggests that fatigue was not a confounding factor in our experiment.
@@ 4 Data Augmentation Experiment
Data augmentation is the process of increasing the amount of data inputted into a machine learning (ML) model by performing the right operations on the data to achieve better results [26, 27, 28, 29]. It helps to counter overfitting, a common issue in ML where the model learns the training data too precisely and fails to generalize.
Commonly used data augmentation techniques for audio ML tasks include pitch shifting, time stretching, speed perturbation, noise addition, harmonic distortion, random time-frequency masking, and amplitude gain [30, 31, 32]. Effect of image-based data augmentation styles on magnitude spectrograms have also been researched [30]. These strategies have improved results for various audio-based tasks.
We propose using phase-intercept distortion as a data augmentation technique for music and speech ML tasks, such as classification, source separation, and generation. This form of data augmentation is a unique way to modify a signal without altering the transients, the pitch, and time characteristics of the signal. Data augmentation can be performed by randomly sampling a $\theta\in[-\pi,\pi)$ and modifying the signal using 7. Since phase-intercept distortion does not affect the magnitude spectrogram (as shown in 3), this augmentation is not applicable to models using magnitude spectrograms as inputs—it is only applicable to those using time-domain signals or complex spectrograms.
To truly test whether the benefits of data augmentation stem from the diversity it provides and not the quantity of data, we designed our experiments with data augmentation added “on the fly” during training. This means that each time a training sample is inputted into the model, it is randomly transformed in real time. This ensures variation without increasing the dataset size, and isolates the impact of diversity provided by data augmentation.
@@ 4.1 Audio Classification
Audio classification is the task of assigning audio signals to corresponding classes based on their characteristics. This process typically involves an initial stage of feature extraction from the raw audio. Some typical features which are extracted include spectrograms, mel-spectrograms, and mel-frequency cepstral coefficients or MFCCs. These features are used as inputs to these classification models that learn to map these features to the audio classes.
Many audio-classification datasets have been published; We chose to use the SC09 dataset, a subset of the Speech Commands dataset [33], which includes spoken digits from zero to nine from various different speakers.
@@
Five different seed values were chosen randomly for the experiment. The mean and standard deviation of the validation and test accuracies are reported in Table 2. The validation accuracy has clearly improved when phase-intercept distortion is used as a data augmentation technique for this model. However, the test accuracy is only slightly better.

### v2

@@ 1.1 Phase-intercept Distortion
$x(t)=\sin(\omega_{0}t+\phi)=\frac{1}{2i}(e^{i\omega_{0}t}e^{i\phi}-e^{-i\omega_{0}t}e^{-i\phi})$ (2)
Note that we must add the positive phase to the positive frequencies and the negative of that phase to the negative frequencies. This can be verified by using a single sine tone and the effect of a shift in phase in its frequency domain.
$\hat{x}(\omega)=\frac{1}{2i}[\delta(\omega-\omega_{0})e^{i\phi}-\delta(\omega+\omega_{0})e^{-i\phi}]$ (3)
@@
A useful property of the Hilbert transform of a real-valued signal $x(t):\mathbb{R}\rightarrow\mathbb{\mathbb{R}}$, is that it is orthogonal to the original signal, as can be easily proved by showing that its inner product with $\mathcal{H}(x(t))$ is equal to zero. This property makes it fundamental in applications such as analytic signal representation [14, 15], envelope detection [16], phase retrieval [17], single-sideband (SSB) modulation, and instantaneous frequency estimation [18].
@@
This can be shown by expanding 6 and rearranging the terms. In summary, phase-intercept distortion can be modeled by constructing the analytic representation of the signal using the Hilbert transform, rotating it, and extracting the real part of this complex signal  [^1] https://github.com/Computational-Cognitive-Musicology-Lab/Phase-Intercept-Distortion.
@@ 1.2 Perception of Phase-intercept Distortion
The perception of phase-intercept distortion has not been explored as much as other common forms of phase distortion. One specific case that has been studied is polarity reversal—corresponding to a phase shift where $\theta=\pm 180^{\circ}$ in 9. Though most researchers (and audio engineers) assume that polarity reversal is inaudible, which is consistent with experimental evidence [19], there have also been some claims that it is audible [7, 8]. The Wood Effect [^2] Named after Charles L. Wood. [8] is one dramatic edge case where polarity reversal creates a minute difference in timbre for lower frequencies. This effect occurs when inverting a sine wave which has been clipped on only one half of each cycle; The resulting polarity-inverted signals sound slightly different, showing that our ears do treat the clipped portion differently when it is presented as a compression or as a rarefaction. No other cases of perceivable phase inversion have been reported.
@@ 2.1 Participants
Forty-eight participants were recruited via different professional and university mailing lists.  [^3] This experiment was conducted with the approval of the Georgia Tech’s Institution Review Board (IRB) (ethics board). Participants were only required to be physically present in the United States, at least 18 years old, and have no history of hearing loss (screened through self-report). Only 27 of our 48 participants completed 100% of the survey and we discarded data from one participant who reported a hearing disability, leaving complete data from 26 participants.
We anticipated that individual differences between participants (age, gender, music background, etc.) might affect their ability to distinguish phase-intercept distortion. Thus, through a pre-questionnaire, we collected information about the participants’ age, gender, and experience with music, musical instruments, music production, mixing engineering, and high-fidelity listening habits. Twenty-two participants identified as male (mean age: 30.4 years) and the rest as female (mean age: 40.8 years); eight of the participants were above the age of 35. Fourteen participants identified as musicians. In the final analysis, none of these factors seem to affect participants’ performance.
@@ 2.3 Audio Editing and Manipulation
We computed the Hilbert transform of each of the thirty stimuli to construct the analytic signals using 8. To finally apply phase-intercept distortion, the sampled $\theta$ values were then input to 9.
@@ 3 Results
For each trial, the participant’s response (A or B) can be coded correct or incorrect. If our hypothesis is true, participants’ true success rate should be approximately 50%. On average, our participants selected correctly in 49.87% of trials (Table 1). A one-sample t-test was performed to check if this accuracy was significantly different from the chance level of 50%, and the result was not significant. However, a non-significant test result might arise from a small sample size, especially if the true effect is small. We thus, focus on Bayesian statistical analysis, which is more appropriate for this analysis.
Table 1: Perceptual Test Results (Percentage of Correct Answers)
@@
Mean 51.15% 45.77% 52.69% 49.87%
Median 50% 50% 50% 50%
Standard deviation 15.27% 19.05% 17.66% 9.36%
t-statistic 0.378 -1.111 0.762 -0.068
p-value 0.709 0.277 0.453 0.946
Samples 26
Figure 3: Distribution of response accuracy of 26 participants.
@@
Out of the 780 total questions answered by the participants, they successfully selected the undistorted stimuli 389 times and failed 391 times. Thus, the Bayesian posterior distribution for $q$, given the naive prior distribution we started with, is $q_{posterior}\sim\textnormal{Beta}(\alpha=1+389,\ \beta=1+391)$ (Fig. 4). As can be seen in Fig. 4, even if we begin with flat prior belief about $q$, the data from this experiment leads to a posterior belief about $q$ that is tightly centered around $q=0.5$, with 95% of this posterior distribution falling between 0.464 and 0.534. These results are consistent with the experimental hypothesis that phase-intercept distortion has no perceptible effects.
@@
Table 1 reports the results for each of the categories: music, speech and other. The median results are 50% for all categories. The mean scores are similar for music and other, while the mean score of speech is worse than 50%.
@@
To ensure that participants’ fatiguing over the course of the survey was not a confounding factor, Fig. 5 plots the question-wise scores, with a fitted trend line. This least-squares regression line has a negligible slope of $-0.002$. This suggests that fatigue was not a confounding factor in our experiment.
@@ 4 Data Augmentation Experiment
In Machine Learning (ML), data augmentation is the process of performing operations to increase the amount of input data to achieve better results [26, 27, 28, 29]. It helps to counter overfitting, a common issue in ML where the model learns the training data too precisely and fails to generalize.
Commonly used data augmentation techniques for audio ML tasks include pitch shifting, time stretching, speed perturbation, noise addition, harmonic distortion, random time-frequency masking, and amplitude gain [30, 31, 32]. The effect of image-based data augmentation styles on magnitude spectrograms have also been researched [30]. These strategies have improved results for various audio-based tasks.
We propose using phase-intercept distortion as a data augmentation technique for music and speech ML tasks, such as classification, source separation, and generation. With an overall computational complexity of $\mathcal{O}(n\log n)$, this form of data augmentation is a fast and unique way to modify a signal without altering the transients, the pitch, and time characteristics of the signal. Data augmentation can be performed by randomly sampling a $\theta\in[-\pi,\pi)$ and modifying the signal using 9. This can be considered as a generalized version of the more restrictive Inverse Phase Augmentation (IPA) [33, 34]. Since phase-intercept distortion does not affect the magnitude spectrogram (as shown in 5), this augmentation is not applicable to models using magnitude spectrograms as inputs—it is only applicable to those using time-domain signals or complex spectrograms.
To truly test whether the benefits of data augmentation stem from the diversity it provides and not the quantity of data, we designed our experiments with data augmentation added “on the fly” during training. This means that each time a training sample is input to the model, it is randomly transformed in real time. This ensures variation without increasing the dataset size, and isolates the impact of diversity provided by data augmentation.
@@ 4.1 Audio Classification
Audio classification is the task of assigning audio signals to corresponding classes based on their characteristics. This process typically involves an initial stage of feature extraction from the raw audio. Some typical features include spectrograms, mel-spectrograms, and mel-frequency cepstral coefficients or MFCCs. These classification models then learn to map these features to the audio classes.
We chose to use the SC09 dataset for the task, a subset of the Speech Commands dataset [35], which includes spoken digits from zero to nine from various different speakers.
@@
Five different seed values were chosen randomly for the experiment. The mean and standard deviation of the validation and test accuracies are reported in Table 2. The validation accuracy shows improvement when phase-intercept distortion is used as a data augmentation technique for this model. However, the test accuracy is only slightly better.